OpenStudy (anonymous):

The length of a rectangle is 5cm more than its width. If its perimeter is 58cm, calculate it dimension and its area

OpenStudy (anonymous):

P = 2L + 2W L = W + 5 P = 58 now we sub 58 = 2(W + 5) + 2W 58 = 2W + 10 + 2W 58 - 10 = 4W 48 = 4W 12 = W L = W + 5 L = 12 + 5 L = 17 Length is 17 cm and width is 12 cm A = L * W A = 17 * 12 A = 204 cm^2 Area is 204 cm^2

OpenStudy (anonymous):

thank you